Find the particular solution of each of the differential equation expressing y explicitly in terms of x.

a) y2dy/dx = 2x2+1 y=1 when x=1

b) xdy/dx = y+2 y=7 when x=3


1
Expert's answer
2021-09-17T03:49:55-0400

a) Integrate


"\int y^2dy=\int (2x^2+1)dx"

"\dfrac{1}{3}y^3 =\dfrac{2}{3}x^3+x+\dfrac{1}{3}C"

"y^3 =2x^3+3x+C"

"y(1)=1"


"(1)^3 =2(1)^3+3(1)+C"

"C=-4"

The particular solution of each of the given differential equation is


"y=\sqrt[3]{2x^3+3x-4}"

b)


"xdy\/dx = y+2"

"\dfrac{dy}{y+2}=\dfrac{dx}{x}"

Integrate


"\int \dfrac{dy}{y+2}=\int\dfrac{dx}{x}"

"\ln(|y+2|)=\ln(|x|)+\ln C"

"y+2=Cx"

"y(3)=7"


"7+2 =C(3)=>C=3"

"C=-4"

The particular solution of each of the given differential equation is


"y=3x-2"